OpenStudy (anonymous):

Good morning, all. Can you please find the real number solution to this equation? p^3-4p^2=0 Much appreciated. x

OpenStudy (akashdeepdeb):

Morning! :D \[p^3 - 4p^2 = 0\] If you see, p^2 is present in both, we can take that common! \[p^2(p - 4) = 0\] So \[p^2 = 0\] and \[p-4 = 0\] Thus the solutions are 0,0,4. I wrote 0 twice because it must have three roots as it is a cubic polynomial. Got it?